OpenStudy (anonymous):

Evaluate ∫ cos2x dx .

OpenStudy (anonymous):

nvm got it

OpenStudy (marissalovescats):

Okay :)

OpenStudy (anonymous):

im just taking college quizes to prepare me for it

OpenStudy (marissalovescats):

If you got 1/2 sin(2x)+C when you're right

OpenStudy (marissalovescats):

then you're right*

OpenStudy (anonymous):

x/2 +(sin 2x)/4 + c is what i got and said i was correct

OpenStudy (marissalovescats):

I believe that's the same answer as mine in a different form

OpenStudy (marissalovescats):

I let u=2x This means du=2dx so 1/2du=dx So 1/2 integral cos u du = 1/2 (sinu) +C 1/2 sin(2x)+C

OpenStudy (anonymous):

Okay thank you.

OpenStudy (marissalovescats):

You're welcome :)